# Fidelity Sweet Spot in Transmon Qubit Rings under Strong Connectivity Noise

### v1

@@ IV.2 Data extraction and neural-network training
To construct the training dataset, we first compute the fidelity as a function of $J/\lambda_{0}$ for system sizes $L=4,6,$ and 8, varying the CPW noise strength within $\lambda_{K}/\lambda_{0}\in[3,20]$ and the noise-distribution width within $\sigma_{K}\in[0.01,0.1]\times\lambda_{K}$. A subset of the resulting data is presented in Appendix D. From these simulations, we extract the sweet-spot locations and their corresponding fidelities to form the target outputs of the machine-learning dataset. The trained neural network is then evaluated using these data, with the resulting predictions shown in Fig. 10.
@@
Figure 10: Comparison between predicted and true values for (a) the sweet-spot location $J^{*}/\lambda_{0}$ and (b) the infidelity $1-\overline{F}^{*}$. The sweet-spot prediction shows excellent agreement, with ${\rm MSE}=0.31$ and $R^{2}\approx 0.99$, whereas fidelity prediction is less accurate, yielding ${\rm MSE}=0.0048$ and $R^{2}\approx 0.81$ on the validation set.
In Fig. 11, we show the evolution of training performance as a function of epoch. Both the MSE and $R^{2}$ metrics converge after approximately 250 epochs, confirming that the chosen neural-network architecture is well suited to the dataset. The prediction of the sweet-spot location is highly accurate, yielding ${\rm MSE}=0.31$ and $R^{2}\approx 0.99$ for the validation set. In comparison, the fidelity prediction performs slightly less well, with ${\rm MSE}=0.0048$ and $R^{2}\approx 0.81$.
@@
Figure 11: Evolution of MSE and coefficient of determination $R^{2}$ during training for (a,b) the sweet-spot position $J^{*}/\lambda_{0}$ and (c,d) the infidelity $1-\overline{F}^{*}$. Both metrics converge after approximately 250 epochs, confirming the stability and effectiveness of the neural-network model.
This difference arises because the sweet-spot location data in the training set are intrinsically more reliable than the fidelity data. As discussed in Sec. III.2, the sweet-spot position depends primarily on system parameters and remains largely independent of the circuit’s specific form, whereas the fidelity is constrained by numerical and computational accuracy. Moreover, the sweet-spot position carries greater experimental relevance: in circuits subject to strong noise or long operation times, there always exists an optimal circuit duration, or equivalently, a circuit depth, that maximizes fidelity, independent of circuit type. Once the parameters are fixed, this optimum is uniquely determined, making its accurate prediction more meaningful than that of the fidelity magnitude. Hence, the performance of our neural-network model fully satisfies the intended purpose.
@@ V Conclusions and discussions
In summary, our investigation of transmon-qubit systems reveals insights into the fidelity of quantum operations. The identification of fidelity sweet spots under realistic noise conditions suggests new routes to enhance gate performance and maintain high operation fidelity in multi-qubit devices. These results are particularly relevant for implementing SWAP and other general operations that underpin robust quantum computing.
We have also examined how different initial states influence operation fidelity, thereby deepening our understanding of quantum-state transport and control. Our results indicate that both the degree of entanglement and the symmetry alignment between the initial state and the quantum circuit play essential roles in determining fidelity, as exemplified by the strong symmetry correspondence between the $|\Psi\rangle_{\mathrm{GHZ}}$ state and the SWAP sequence. The influence of multi-qubit state symmetries on noise resilience remains an open question that merits further theoretical and experimental study.
Finally, the integration of machine-learning techniques provides a powerful and resource-efficient means of predicting fidelity sweet spots and optimizing circuit durations without exhaustive simulations or measurements. Our supervised neural-network approach adapts effectively to variations in CPW coupling strength and noise distributions, enabling rapid identification of optimal operation parameters across diverse devices. Overall, we hope that the results presented in this work may offer some useful insight for the design of next-generation, high-fidelity quantum processors.
@@ Acknowledgements.
The authors thank Liu Jie, Wu Jiaohao and Ni Xiaotong for fruitful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12474489), the Shenzhen Fundamental Research Program (Grant No. JCYJ20240813153139050), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2403001), and Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2021ZD0302300).
DATA AVAILABILITY
The data that support the findings of this article are openly available Fu and Wang (2025).
@@ Appendix A Comprehensive Model for Transmon Qubits with Hybrid Connectivity
We present a comprehensive Hamiltonian model for systems of L=4L=4,6, and 8 transmon qubits arranged in a ring geometry with hybrid cavity-CPW connectivity, as illustrated in Fig. 1 Koch et al. (2007). The model includes both static and dynamic noise contributions that persist independently of the applied control operations, indicating that the noise characteristics are determined primarily by intrinsic device parameters and fabrication quality Zhang et al. (2019); Zhao et al. (2025a).
@@ A.2.2 CPW photon loss-induced noise
The scaling $\lambda_{ij,0}^{K}\propto\sqrt{\kappa}$ highlights the fundamental link between photon loss and noise amplitude: increased photon dissipation amplifies vacuum fluctuations, thereby elevating the noise floor. Similarly, the variance $\sigma_{K}^{2}$ follows a comparable dependence on photon loss and mode density,
$\sigma_{K}\propto\frac{\bar{g}_{i}\bar{g}_{j}\kappa^{3/4}}{\bar{\omega}^{1/2}}$ (47)
confirming the intrinsic connection between CPW photon loss ($\kappa=\bar{\omega}/Q$) and the statistical properties of coupling noise.
@@ A.3 Complete system Hamiltonian
Table 1: Experimental parameters for transmon devices Kong (2018)
@@
Parasitic capacitance noise, $\lambda^{c}_{ij}(t)$ 0.05–0.5 MHz
Photon loss noise, $\lambda^{t}_{ij}(t)$ 1.0–5.0 MHz
@@ Appendix B Generating Random Unitary Operations
Appendix C Analytical origin of the fidelity sweet spot: a two-qubit example
To elucidate the physical origin of the fidelity sweet spot, we analyze a minimal two-qubit model undergoing two consecutive SWAP operations. Although such a system does not require both cavity and CPW couplings, we employ a simplified Hamiltonian to extract analytical insight into coherence behavior. The model Hamiltonian is defined as
@@ Appendix C Analytical origin of the fidelity sweet spot: a two-qubit example
Figure 13: Fidelity of two-qubit states under Gaussian noise during two SWAP operations for different $\sigma$. The $|\uparrow\downarrow\rangle$ state shows a clear sweet spot at small $J/\lambda_{0}$, consistent with numerical simulations. In contrast, the $|\text{GHZ}\rangle$ (Bell) state exhibits no sweet spot because the simplified model omits qubit frequency terms, rendering it a Hamiltonian eigenstate and thus suppressing oscillatory evolution.
@@
These expressions explicitly show how the interplay between coupling strength and Gaussian noise determines the fidelity of both quantum states. By setting $\lambda_{ij,0}^{t}=10,\lambda_{ij,0}^{c}=\lambda_{0}$, we take $a=10/11$. As shown in Fig. 13, the state $\left|\uparrow\downarrow\right\rangle$ exhibits a distinct sweet spot at small $J/\lambda_{0}$, consistent with the numerical results presented in the main text. In contrast, the two-qubit $|\text{GHZ}\rangle$ (Bell) state shows no visible sweet spot because, in this simplified toy model, the omission of superconducting-qubit frequency terms makes $|\text{GHZ}\rangle$ an eigenstate of the Hamiltonian, thereby suppressing oscillatory evolution. When these frequency terms are restored, the $|\text{GHZ}\rangle$ state is expected to exhibit an analogous sweet spot behavior.
@@ Appendix D Sweet-spot dependence on interaction strength
This variation physically corresponds to modifying the CPW parameters, with each configuration leading to a unique optimal operation time associated with its sweet spot. Once the optimal parameter is identified, other unitary operations in the same device can be implemented with comparably high fidelity, consistent with the findings of Sec. III.2. Interestingly, increasing $\lambda^{K}$ lowers the sweet-spot position (yielding higher fidelity) but also shortens the gate time. This result highlights a practical trade-off between operation duration and achievable fidelity, providing experimentally relevant guidance for system optimization.
@@ Appendix E Time-dependent noise evolution
The fidelity between the ideal and noisy evolutions is then evaluated as
$F=\mathrm{Tr}\left[\left(\prod_{k=1}^{K}U_{k}\right)\rho_{0}\left(\prod_{k=1}^{K}U_{k}\right)^{\dagger}\rho_{\mathrm{final}}\right].$ (75)

### v2

@@ IV.2 Data extraction and neural network training
To construct the training dataset, we compute fidelity as a function of $J/\lambda_{0}$ for system sizes $L=4,6,$ and 8, varying CPW noise strength within $\lambda^{K}/\lambda_{0}\in[3,20]$ and noise-distribution width within $\sigma_{K}\in[0.01,0.1]\times\lambda^{K}$. From these simulations, we extract optimal operation point positions and corresponding fidelities to form the target outputs.
Figure 10 shows the resulting predictions. The optimal point position prediction shows excellent agreement with validation data (${\rm MSE}=0.31$, $R^{2}\approx 0.99$), while fidelity prediction is less accurate (${\rm MSE}=0.0048$, $R^{2}\approx 0.81$). This difference arises because optimal point positions are more reliably determined from simulations and exhibit greater consistency across different circuits, as discussed in Section III.2.
@@
Figure 10: Comparison between predicted and true values for (a) optimal operation point position $J^{*}/\lambda_{0}$ and (b) infidelity $1-\overline{F}^{*}$. Position prediction shows excellent agreement (${\rm MSE}=0.31$, $R^{2}\approx 0.99$), while fidelity prediction is less accurate (${\rm MSE}=0.0048$, $R^{2}\approx 0.81$).
Figure 11 shows the evolution of training performance. Both MSE and $R^{2}$ metrics converge after approximately 250 epochs, confirming the stability and effectiveness of the neural network model. The accurate prediction of optimal operation point positions is particularly valuable experimentally, as it enables determination of ideal circuit durations without extensive characterization measurements.
@@
Figure 11: Evolution of MSE and $R^{2}$ during training for (a,b) optimal point position $J^{*}/\lambda_{0}$ and (c,d) infidelity $1-\overline{F}^{*}$. Both metrics converge after approximately 250 epochs.
@@ V Conclusions and discussions
In summary, our investigation of transmon qubit systems reveals that optimal operation points emerge consistently across various quantum operations, even under substantial connectivity noise. The identification of these fidelity maxima in the intermediate coupling regime ($10\lesssim J/\lambda_{0}\lesssim 100$) provides a practical pathway to enhance gate performance in multiqubit devices without requiring unrealistic noise suppression. This finding is particularly relevant for implementing SWAP gates and general quantum operations that form the foundation of scalable quantum computing architectures.
Our analysis demonstrates that the position of optimal operation points remains largely independent of the specific quantum circuit, enabling efficient optimization across diverse operations. Moreover, the significant dependence of fidelity on initial states reveals that entanglement structure and symmetry alignment play crucial roles in noise resilience. The exceptional performance of GHZ states under SWAP operations, reaching fidelities compatible with quantum error correction thresholds, highlights the importance of state-circuit symmetry matching for robust quantum information processing. The underlying mechanism, rooted in coherence oscillations under quasistatic noise, provides a universal framework for understanding fidelity dynamics in superconducting qubit systems.
The integration of machine learning techniques further enhances the practical utility of our findings by enabling efficient prediction of optimal operation points across varying device parameters. Our supervised neural network approach accurately identifies optimal circuit durations without exhaustive simulations, adapting effectively to variations in CPW coupling strengths and noise distributions. This data driven strategy significantly reduces the experimental overhead for optimizing quantum operations in diverse device configurations.
Overall, our results provide valuable insights for designing next generation quantum processors that leverage optimal gate duration tuning to achieve high fidelity operations under realistic noise conditions. By identifying and exploiting the inherent trade-offs between circuit depth and noise accumulation, these strategies offer a pragmatic approach to enhancing quantum computational performance in the noisy intermediate scale quantum era.
@@ Acknowledgements.
The authors thank Liu Jie, Wu Jiaohao, He Minquan, Zhu Yan and Ni Xiaotong for fruitful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12474489), the Shenzhen Fundamental Research Program (Grant No. JCYJ20240813153139050), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2403001), and Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2021ZD0302300).
Declarations
The authors declare no conflict of interest.
@@ Appendix A Comprehensive Model for Transmon Qubits with Hybrid Connectivity
We present a comprehensive Hamiltonian model for systems of L=4L=4,6, and 8 transmon qubits arranged in a ring geometry with hybrid cavity-CPW connectivity, as illustrated in Fig. 1 Koch et al. (2007). The model includes both static and dynamic noise contributions that persist independently of the applied control operations, indicating that the noise characteristics are determined primarily by intrinsic device parameters and fabrication quality Zhang et al. (2019); Zhao et al. (2025b).
@@ A.1 Control Hamiltonian
We note that the coupling strength $J_{ij}$ depends on the detunings $\Delta_{ik}$ and $\Delta_{jk}$. In experimental systems, tuning the frequency of a qubit (thus changing its detuning from a resonator) will affect all coupling terms involving that qubit. For instance, varying $\Delta_{ik}$ to adjust $J_{ij}$ will also modify $J_{i\ell}$ for other qubits $\ell$ coupled to the same resonator $k$. This interdependence implies that in practice, calibrating multiple couplings may require global optimization over several parameters. In our theoretical model, we treat the $J_{ij}(t)$ as independently tunable parameters, which is an idealization that allows us to focus on the generic effects of connectivity noise. Advanced control techniques, such as using dedicated tunable couplers, can provide a high degree of independent control over effective couplings, approximating this idealization. Similar considerations apply to the CPW-mediated couplings $K_{ij}$.
@@ A.2.2 CPW photon loss-induced noise
The scaling $\lambda_{ij,0}^{K}\propto\sqrt{\kappa}$ highlights the fundamental link between photon loss and noise amplitude: increased photon dissipation amplifies vacuum fluctuations, thereby elevating the noise floor. Similarly, the variance $\sigma_{K}^{2}$ follows a comparable dependence on photon loss and mode density, $\sigma_{K}\propto\frac{\bar{g}_{i}\bar{g}_{j}\kappa^{3/4}}{\bar{\omega}^{1/2}}$, confirming the intrinsic connection between CPW photon loss ($\kappa=\bar{\omega}/Q$) and the statistical properties of coupling noise.
The effective noise amplitude $\lambda_{ij}^{K}(t)$ scales as $\propto 1/(\Delta_{ik}\Delta_{jk})$. In practical devices, tuning the qubit frequency (thus changing $\Delta_{ik}$) to adjust the intended coupling $J_{ij}$ would simultaneously modify the noise statistics. In our theoretical model, however, we treat the noise amplitudes $\lambda_{ij}^{J}$ and $\lambda_{ij}^{K}$ as fixed parameters determined by static device properties. This simplification allows us to focus on the generic dependence of fidelity on the dimensionless ratio $J/\lambda_{0}$. We emphasize that the existence of optimal operation points is a robust phenomenon that persists even when noise amplitudes vary with detuning, as the underlying physical mechanism is governed by the ratio between coupling and noise energy scales. For experimental calibration, a global optimization over both coupling and noise parameters may be required.
@@ A.3 Complete system Hamiltonian
Table 1: Simulation parameters for transmon devices, based on typical experimental ranges Kong (2018); Wang et al. (2022). The chosen values allow isolation of connectivity-noise effects; the results exhibit scaling invariance as discussed in the text.
@@
Parasitic capacitance noise, $\lambda^{J}_{ij}(t)$ 0.05–0.5 MHz
Photon loss noise, $\lambda^{K}_{ij}(t)$ 1.0–5.0 MHz
Noise ratio, $\mathcal{R}=\lambda^{K}_{ij}/\lambda^{J}_{ij}$ 3–20
@@ Appendix B Generating Random Unitary Operations
Appendix C Analytical origin of the fidelity optimal operation point: a two qubit example
To elucidate the physical origin of the fidelity optimal operation point, we analyze a minimal two-qubit model undergoing two consecutive SWAP operations. Although such a system does not require both cavity and CPW couplings, we employ a simplified Hamiltonian to extract analytical insight into coherence behavior. The model Hamiltonian is defined as
@@ Appendix C Analytical origin of the fidelity optimal operation point: a two qubit example
Figure 13: Fidelity of two-qubit states under Gaussian noise during two SWAP operations for different $\sigma$. The $|\uparrow\downarrow\rangle$ state shows a clear optimal operation point at small $J/\lambda_{0}$, consistent with numerical simulations. In contrast, the $|\text{GHZ}\rangle$ (Bell) state exhibits no optimal operation point because the simplified model omits qubit frequency terms, rendering it a Hamiltonian eigenstate and thus suppressing oscillatory evolution.
@@
These expressions explicitly show how the interplay between coupling strength and Gaussian noise determines the fidelity of both quantum states. By setting $\lambda_{ij,0}^{t}=10,\lambda_{ij,0}^{c}=\lambda_{0}$, we take $a=10/11$. As shown in Fig. 13, the state $\left|\uparrow\downarrow\right\rangle$ exhibits a distinct optimal operation point at small $J/\lambda_{0}$, consistent with the numerical results presented in the main text. In contrast, the two-qubit $|\text{GHZ}\rangle$ (Bell) state shows no visible optimal operation point because, in this simplified toy model, the omission of superconducting-qubit frequency terms makes $|\text{GHZ}\rangle$ an eigenstate of the Hamiltonian, thereby suppressing oscillatory evolution. When these frequency terms are restored, the $|\text{GHZ}\rangle$ state is expected to exhibit an analogous optimal operation point behavior.
@@ Appendix D Sweet-spot dependence on interaction strength
This variation physically corresponds to modifying the CPW parameters, with each configuration leading to a unique optimal operation time associated with its optimal operation point. Once the optimal parameter is identified, other unitary operations in the same device can be implemented with comparably high fidelity, consistent with the findings of Sec. III.2. Interestingly, increasing $\lambda^{K}$ lowers the sweet-spot position (yielding higher fidelity) but also shortens the gate time. This result highlights a practical trade-off between operation duration and achievable fidelity, providing experimentally relevant guidance for system optimization.
@@ Appendix E Time-dependent noise evolution
The fidelity reported throughout the manuscript is computed as the overlap between the final state under noisy evolution and the ideal (noise-free) final state. For a given initial state $|\psi_{0}\rangle$, we first compute the ideal final state $|\psi_{\text{ideal}}\rangle=U_{\text{ideal}}|\psi_{0}\rangle$, where $U_{\text{ideal}}$ is the target unitary (e.g., the product of SWAP gates or a general random unitary). The noisy evolution yields the final density matrix $\rho_{\text{final}}$ via the above Trotter procedure. The fidelity is then given by
$F=\langle\psi_{\text{ideal}}|\rho_{\text{final}}|\psi_{\text{ideal}}\rangle.$ (74)
If the noisy evolution is pure (i.e., $\rho_{\text{final}}=|\psi_{\text{actual}}\rangle\langle\psi_{\text{actual}}|$), this reduces to $F=|\langle\psi_{\text{ideal}}|\psi_{\text{actual}}\rangle|^{2}$. This definition coincides with the standard quantum state transfer (QST) fidelity when the target operation is a state transfer protocol. In our SWAP circuits, the ideal operation cyclically permutes the qubit states so that the ideal output equals the input; thus the fidelity measures the success probability of returning the initial state, which is exactly the QST fidelity. For general random circuits, the fidelity quantifies how accurately the noisy implementation reproduces the target unitary operation, generalizing the QST fidelity to arbitrary multi-qubit transformations.